# Probing cosmic isotropy with Gamma-ray bursts: A dipole and quadrupole analysis of BATSE and Fermi GBM data

### v1

@@ Abstract
The cosmological principle, asserting large-scale homogeneity and isotropy, underpins the standard model of cosmology. Testing its validity using independent astronomical probes remains crucial for understanding the global structure of the Universe. We investigate the angular distribution of Gamma-Ray Bursts (GRBs) using two of the most comprehensive all-sky datasets available, the BATSE (CGRO) and Fermi GBM catalog to test the isotropy of the GRB sky at large angular scales. We perform spherical harmonic decomposition of the GRB sky maps and estimate the dipole and quadrupole amplitudes. Statistical significance is evaluated by comparing the observed multipole amplitudes against distributions derived from 500 Monte Carlo realizations of isotropic skies. Our results show that the observed dipole amplitudes for both BATSE and Fermi GBM datasets lie within the $1\sigma$ region of their respective null distributions. However, the quadrupole amplitude in the raw BATSE and Fermi skies appears elevated at $3.7\sigma$ and $3.0\sigma$ respectively. After incorporating the BATSE sky exposure function, this apparent quadrupole anisotropy vanishes, indicating that instrumental non-uniformities fully account for the signal. Our method’s reliability is validated through controlled simulations, which show it can detect the injected dipoles in BATSE-sized isotropic skies. These findings reinforce the statistical isotropy of the GRB sky and underscore the importance of accurate exposure corrections in cosmological anisotropy analyses.
@@ 1 Introduction
A non-zero dipole in the distribution of GRBs would indicate a directional anisotropy, a statistically significant excess of GRBs in one region of the sky compared to its opposite. Similarly, a non-zero quadrupole in the distribution of GRBs would indicate a planar or axial anisotropy, suggesting a systematic variation in the number of bursts across two opposite hemispheres or along a preferred plane in the sky. These would challenge the cosmological principle, which assumes the Universe is isotropic on large scales. In this study, we examine the angular isotropy of gamma-ray bursts using a statistical framework applied to two of the most comprehensive and widely used GRB datasets: the BATSE and Fermi GBM catalogs. We measure the dipoles and quadrupoles in these datasets and assess their statistical significance. While previous analyses have considered dipole patterns in GRB distributions (Briggs et al. 1996), our methodology advances this line of inquiry by generating empirical dipole amplitude distributions from Monte Carlo simulations of isotropic skies. By comparing the observed dipole amplitudes to these simulated distributions, we obtain a direct and statistically robust estimate of the significance of any detected anisotropy. To further validate the reliability of our approach, we perform control tests using mock GRB skies with injected dipole signatures, thereby ensuring that our framework is sensitive to known deviations from isotropy. This simulation-driven methodology provides a systematic and data-driven means of testing the null hypothesis of angular isotropy in GRB distributions.
@@ 2.2.4 Validation with simulated dipole skies
The results clearly demonstrate that increasing the number of data points enhances the sensitivity of the analysis. In this higher-resolution test, even a weak dipole signal with amplitude $a=0.022$ lies beyond the $2\sigma$ region of the isotropic distribution, confirming that the method can reliably detect subtle anisotropies when applied to sufficiently large datasets. This underscores that the efficiency of dipole detection depends not only on the signal strength but also on the statistical size of the sample representing the GRB sky.
@@ 3.4 Quadrupole amplitude analysis of the FERMI dataset
Although the BATSE dataset allows for a rigorous treatment of instrumental exposure through the availability of its well-characterized 5B exposure function, no equivalent full-sky exposure maps are publicly available for the Fermi GBM instrument. Consequently, we were unable to perform a similar exposure-corrected isotropy analysis for the Fermi GRB dataset. This limitation is primarily due to the complex observational geometry of Fermi’s sky coverage, which varies with both time and energy threshold as the satellite operates in an all-sky survey mode. Nevertheless, the results obtained from the BATSE analysis provide a strong indication of the likely origin of the apparent anisotropies seen in the Fermi data. Given that the Fermi GBM exhibits a non-uniform sky exposure pattern driven by its orbital constraints, Earth occultation, and South Atlantic Anomaly passages, it is reasonable to expect that the observed $3\sigma$ quadrupole signal in the Fermi GRB distribution may also arise from residual exposure-related systematics rather than a genuine cosmological anisotropy. A proper exposure correction would likely reduce or even eliminate this apparent excess, as seen for the BATSE case. Therefore, while the Fermi results remain intriguing, they should be interpreted with caution until a detailed, exposure-weighted isotropy test becomes feasible with future data releases or reconstructed exposure models.
@@ 4 Conclusions
To investigate the impact of instrumental effects, we further applied exposure corrections to the BATSE data using its publicly available 5B exposure map. After accounting for non-uniform sky exposure, both the dipole and quadrupole signals are found to be statistically consistent with isotropy, suggesting that the earlier quadrupole excess was primarily instrumental in origin. Unfortunately, due to the absence of a publicly available full-sky exposure model for Fermi GBM, a similar correction could not be applied to that dataset. Nevertheless, the analogy with BATSE strongly suggests that the apparent Fermi quadrupole anisotropy is also likely to be an artifact of non-uniform sky exposure rather than a cosmological signal.
@@
Our findings are in broad agreement with those that report no significant evidence for large-scale anisotropy in GRB distributions (Briggs et al. 1996; Bernui et al. 2008; Řípa and Shafieloo 2019; Andrade et al. 2019). In particular, our results align with works that found long-duration GRBs to be consistent with isotropy, and contribute additional support using an updated, simulation-backed dipole analysis. At the same time, our conclusions contrast with some earlier claims of anisotropy or clustering in GRB datasets (Mészáros 2019; Vavrek et al. 2008; Balázs et al. 1999; Tarnopolski 2017), particularly among short-duration GRBs. These discrepancies highlight the importance of distinguishing between GRB subpopulations and ensuring uniform analysis procedures. Unlike some previous efforts, our method is validated through controlled simulations with known dipole injections, demonstrating its sensitivity to even weak dipole signals.
Overall, our findings reinforce the validity of the cosmological principle on the largest observable scales probed by GRBs. The lack of statistically significant dipole or quadrupole anisotropies once exposure effects are properly accounted for indicates that GRBs trace an isotropic Universe, consistent with the expectations of the standard cosmological model. Future missions with improved sky coverage, uniform exposure, and larger GRB samples will allow for even more stringent tests, potentially revealing subtler deviations that could provide deeper insights into the large-scale structure and evolution of the universe.
@@ Appendix A Appendix: Stability tests with different HEALPix resolutions
A.1 Effects of HEALPix resolution parameter on dipole and quadrupole estimates

### v2

@@ Abstract
The cosmological principle, asserting large-scale homogeneity and isotropy, underpins the standard model of cosmology. Testing its validity using independent astronomical probes remains crucial for understanding the global structure of the Universe. We investigate the angular distribution of Gamma-Ray Bursts (GRBs) using two of the most comprehensive all-sky datasets available, the BATSE (CGRO) and Fermi GBM catalogs, to test the isotropy of the GRB sky at large angular scales. We perform spherical harmonic decomposition of the GRB sky maps and estimate the dipole and quadrupole amplitudes. Statistical significance is evaluated by comparing the observed multipole amplitudes against distributions derived from 500 Monte Carlo realizations of isotropic skies. Our results show that the observed dipole amplitudes for both BATSE and Fermi GBM datasets lie within the $1\sigma$ region of their respective null distributions. However, the quadrupole amplitude in the raw, uncorrected BATSE and Fermi GBM skies appears elevated at $3.7\sigma$ and $3.0\sigma$, respectively. After incorporating the BATSE sky exposure function, this apparent quadrupole anisotropy vanishes, indicating that instrumental non-uniformities fully account for the signal in that case. Owing to the absence of a publicly available full-sky exposure model for Fermi GBM, the Fermi analysis is restricted to the raw sky distribution. Our method’s reliability is validated through controlled simulations, which show it can detect the injected dipoles in BATSE-sized isotropic skies. These findings reinforce the statistical isotropy of the GRB sky and underscore the importance of accurate exposure corrections in cosmological anisotropy analyses.
@@ 1 Introduction
A non-zero dipole in the distribution of GRBs would indicate a directional anisotropy, a statistically significant excess of GRBs in one region of the sky compared to its opposite. Similarly, a non-zero quadrupole in the distribution of GRBs would indicate a planar or axial anisotropy, suggesting a systematic variation in the number of bursts across two opposite hemispheres or along a preferred plane in the sky. These would challenge the cosmological principle, which assumes the Universe is isotropic on large scales.
In this study, we examine the angular isotropy of gamma-ray bursts using a statistical framework applied to two of the most comprehensive and widely used GRB datasets: the BATSE and Fermi GBM catalogs. We consider the combined GRB sample without separating long- and short-duration bursts. This approach implicitly assumes that, at the largest angular scales probed here, the combined GRB population provides a meaningful tracer of the underlying sky distribution. However, the long-duration and short-duration GRBs arise from different astrophysical progenitors, tracing different cosmic environments and evolutionary histories. The implications of this assumption for the interpretation of our results are discussed in the conclusions section. We measure the dipoles and quadrupoles in these datasets and assess their statistical significance. While previous analyses have considered dipole patterns in GRB distributions (Briggs et al. 1996), our methodology advances this line of inquiry by generating empirical dipole amplitude distributions from Monte Carlo simulations of isotropic skies. By comparing the observed dipole amplitudes to these simulated distributions, we obtain a direct and statistically robust estimate of the significance of any detected anisotropy. To further validate the reliability of our approach, we perform control tests using mock GRB skies with injected dipole signatures, thereby ensuring that our framework is sensitive to known deviations from isotropy. This simulation-driven methodology provides a systematic and data-driven means of testing the null hypothesis of angular isotropy in GRB distributions.
@@ 2.2.4 Validation with simulated dipole skies
The results clearly demonstrate that increasing the number of data points enhances the sensitivity of the analysis. In this higher-resolution test, even a weak dipole signal with amplitude $a=0.022$ lies beyond the $2\sigma$ region of the isotropic distribution, confirming that the method can reliably detect subtle anisotropies when applied to sufficiently large datasets. This underscores that the efficiency of dipole detection depends not only on the signal strength but also on the statistical size of the sample representing the GRB sky. Larger simulated samples would allow the detection of progressively weaker dipole signals at a fixed confidence level. The choice of $20000$ points in our validation tests is intended to illustrate the behavior of the method in a high-statistics regime, beyond the sizes of the current BATSE and Fermi GBM catalogs. Using a smaller number of events (e.g., $10000$) would require a correspondingly larger dipole amplitude to achieve the same level of statistical significance, while larger samples (e.g., $50000$ points) would enable sensitivity to even weaker anisotropies. Varying the number of simulated points does not alter the qualitative performance of the method, but instead produces the expected scaling of detection thresholds with sample size.
@@ 3.4 Quadrupole amplitude analysis of the FERMI dataset
Motivated by the exposure-corrected BATSE results discussed above, we now revisit the interpretation of the Fermi GBM quadrupole signal. For the Fermi GBM dataset, we find that while the dipole amplitude remains fully consistent with isotropy, the quadrupole amplitude exhibits a deviation at the $\sim 3\sigma$ level when compared with isotropic simulations. Although such a level of statistical significance is not sufficient to establish a definitive departure from isotropy, it is commonly regarded as suggestive in cosmological analyses and therefore merits careful consideration. The BATSE dataset allows for a rigorous treatment of instrumental exposure through the availability of its well-characterized 5B exposure function, but no equivalent full-sky exposure maps are publicly available for the Fermi GBM instrument. Consequently, we were unable to perform a similar exposure-corrected isotropy analysis for the Fermi GRB dataset. The physical interpretation of the quadrupole excess in the Fermi data must therefore be treated with caution.
In addition to instrumental exposure non-uniformities, several other observational effects may plausibly contribute to the quadrupole feature observed in the Fermi GBM GRB sky. The GBM instrument employs multiple detectors with different orientations and energy-dependent sensitivities, and its effective sky coverage varies with time due to orbital precession, Earth occultation, and passages through the South Atlantic Anomaly. These factors can introduce large-scale modulations in detection efficiency that are not strictly isotropic and may project preferentially onto low-order multipoles such as the quadrupole. Furthermore, the GBM trigger algorithm operates with background-dependent thresholds that vary with detector orientation, geomagnetic environment, and sky position, potentially introducing subtle selection biases. Even modest, percent-level variations in detection efficiency across broad sky regions could give rise to apparent quadrupole amplitudes of the magnitude observed here. In the absence of detailed, time-resolved exposure and trigger-efficiency maps for Fermi GBM, it is not currently possible to disentangle the relative contributions of these effects from a potential intrinsic signal. These considerations reinforce the interpretation of the Fermi quadrupole excess as a moderate-significance and potentially instrumental feature, warranting caution and motivating future analyses based on improved exposure characterization.
Although the analogy with the BATSE analysis where the apparent quadrupole anisotropy vanishes after accounting for instrumental exposure suggests that the Fermi quadrupole signal is likely influenced by residual exposure-related systematics, we cannot definitively rule out the presence of a weak intrinsic quadrupole component in the Fermi GRB distribution. Furthermore, since the observed Fermi dipole amplitude lies slightly below the mean of the isotropic distribution, the effect of a hypothetical exposure correction on the dipole and quadrupole significances cannot be reliably anticipated. The Fermi results thus remain intriguing, and the observed quadrupole excess should be regarded as a moderate-significance feature and a potential anomaly rather than conclusively attributed to either instrumental or cosmological origins.
@@ 4 Conclusions
To investigate the impact of instrumental effects, we further applied exposure corrections to the BATSE data using its publicly available 5B exposure map. After accounting for non-uniform sky exposure, both the dipole and quadrupole signals are found to be statistically consistent with isotropy, suggesting that the earlier quadrupole excess was primarily instrumental in origin. Unfortunately, due to the absence of a publicly available full-sky exposure model for Fermi GBM, a similar correction could not be applied to that dataset. The analogy with BATSE suggests that the apparent Fermi quadrupole anisotropy is also likely to be an artifact of non-uniform sky exposure rather than a cosmological signal. However, the apparent quadrupole excess in the Fermi GBM dataset should be interpreted cautiously in the absence of an exposure-corrected analysis. Future GRB datasets with improved exposure characterization and redshift completeness will be essential for determining whether this feature persists as a genuine cosmological signal or is fully explained by observational effects.
When interpreting the statistical significances reported above, it is also important to note that they are primarily based on 500 Monte Carlo realizations of isotropic GRB skies. With this number of simulations, the minimum directly resolvable tail probability is of order $p\sim 2\times 10^{-3}$ (Table 1), implying that significance estimates approaching the $\sim 3\sigma$ level are subject to additional uncertainty arising from finite sampling of the distribution tails. To mitigate this limitation, we construct smooth probability density functions using kernel density estimation (KDE), which reduces the impact of shot noise and sampling fluctuations in the extreme tails of the empirical distributions. We therefore adopt a deliberately conservative interpretation of results near the $\sim 3\sigma$ level, treating them as moderate-significance and suggestive rather than definitive detections. Importantly, our main conclusions such as the disappearance of the BATSE quadrupole signal after exposure correction and the consistency of dipole amplitudes with isotropy do not rely on precise tail probabilities and remain robust against modest variations in the estimated significance. Larger simulation ensembles would be required for higher-precision tail characterization, but are not expected to qualitatively alter the conclusions presented here.
@@
Our findings are in broad agreement with those that report no significant evidence for large-scale anisotropy in GRB distributions (Briggs et al. 1996; Bernui et al. 2008; Řípa and Shafieloo 2019; Andrade et al. 2019). In particular, our results align with works that found long-duration GRBs to be consistent with isotropy, and contribute additional support using an updated, simulation-backed dipole analysis. At the same time, our conclusions contrast with some earlier claims of anisotropy or clustering in GRB datasets (Mészáros 2019; Vavrek et al. 2008; Balázs et al. 1999; Tarnopolski 2017), particularly among short-duration GRBs. These discrepancies highlight the importance of distinguishing between GRB subpopulations. Given their distinct astrophysical origins and environments, these two populations may, in principle, exhibit different large-scale anisotropy signatures, as suggested by several earlier studies. If the long- and short-duration GRB populations possess anisotropies of differing amplitudes or orientations, a combined analysis could partially suppress or mask such signals, yielding an apparently isotropic distribution. While separating the sample into sub-classes would provide a more physically refined test, the significantly reduced sample sizes particularly for short-duration GRBs and additional classification and selection uncertainties currently limit the statistical power of such an approach. Our results should therefore be interpreted as constraints on the isotropy of the combined GRB population, and future studies with larger, better-classified samples will be essential for probing anisotropy in individual GRB sub-populations.
Further, several studies have reported large-scale spatial features in the GRB distribution, such as the Giant GRB Ring and other clustering signals identified using redshift information and three-dimensional spatial analyses. It is important to clarify that these results probe a different regime of the GRB distribution than the present work. Our analysis is specifically designed to test large-scale angular isotropy through low-order spherical harmonic moments, which are sensitive to global, coherent anisotropies across the entire sky. In contrast, reported GRB structures are typically identified using localized clustering or pattern-recognition techniques in three-dimensional space and may be confined to specific redshift ranges or angular regions. Consequently, the angular scales and statistical methodologies involved are not directly comparable. The absence of significant dipole anisotropy and the consistency with isotropy at the quadrupole level once instrumental exposure effects are accounted for do not rule out the existence of such localized or redshift-dependent GRB structures. Rather, our results place constraints on the presence of large-scale, all-sky anisotropies in the GRB distribution, implying that any previously reported structures, if real, do not dominate the global angular distribution of GRBs or produce a detectable imprint in the lowest angular multipoles.
We also emphasize that the present study is restricted to the two-dimensional angular distribution of GRBs on the sky. While GRBs are intrinsically three-dimensional cosmological tracers, a fully spatial isotropy test or an analysis in redshift shells (da Silveira Ferreira and Marra 2024) remains challenging with current data. Only a relatively small fraction of GRBs detected by all-sky instruments such as BATSE and Fermi GBM have reliably measured redshifts, and these subsamples are affected by strong selection effects arising from follow-up strategies, detector sensitivities, and sky coverage. Under these circumstances, angular analyses provide the most statistically robust and least biased framework for testing large-scale isotropy. The results presented here should therefore be regarded as a necessary and well-motivated first step. Future GRB catalogs with substantially larger and more homogeneous redshift coverage will be essential for extending these tests into three dimensions and for probing the isotropy of the Universe as a function of cosmic epoch.
Despite these limitations, the consistency of our results across two independent GRB catalogs, spanning different missions and observational epochs, provides strong support for the statistical isotropy of the Universe on the largest angular scales accessible with current GRB data. Our analysis demonstrates that, once instrumental exposure effects are properly accounted for, neither dipole nor quadrupole anisotropies persist in the BATSE GRB sky, and that the Fermi GBM results are consistent with this picture within present observational constraints. As GRB samples continue to grow and redshift completeness improves through sustained multi-wavelength follow-up and future high-energy missions, the methodology presented here can be naturally extended to redshift-resolved analyses. Such developments will enable increasingly stringent tests of the Cosmological Principle and further clarify the role of GRBs as powerful probes of the large-scale structure and evolution of the Universe.
@@ Appendix A Appendix: Stability tests with different HEALPix resolutions
A.1 Effects of HEALPix resolution parameter on dipole and quadrupole estimates for the exposure-corrected BATSE data
@@
A.2 Stability of dipole and quadrupole estimates for raw BATSE and Fermi GRB data
In addition to the stability tests performed for the exposure-corrected BATSE dataset, we also examine the robustness of the dipole and quadrupole estimates obtained from the raw, uncorrected GRB sky maps. This analysis is motivated by the need to verify that the elevated quadrupole amplitudes identified in the raw BATSE and Fermi datasets are not artifacts of the chosen HEALPix pixelization scale, but instead represent stable features of the uncorrected sky distributions.
Following the same procedure adopted in the main analysis, we recompute the dipole and quadrupole amplitudes for the raw BATSE and Fermi GRB skies using higher HEALPix resolutions, specifically $N_{\mathrm{side}}=32$ and $N_{\mathrm{side}}=64$, in addition to the baseline resolution $N_{\mathrm{side}}=8$. For each resolution, the observed multipole amplitudes are compared with probability density functions constructed from 500 Monte Carlo realizations of isotropic skies containing the same number of sources as the corresponding datasets.
For the raw BATSE data, the quadrupole amplitude remains consistently elevated relative to the isotropic expectation across all tested resolutions, with the observed deviation remaining close to the $\sim 3.9\sigma$ level (Figures 20 and 21). Similarly, the dipole amplitude continues to lie well within the isotropic distribution, independent of the chosen pixelization scale (Figures 18 and 19). This confirms that the apparent quadrupole excess observed in the raw BATSE sky is not driven by the choice of $N_{\mathrm{side}}$, but reflects a stable property of the uncorrected map.
An analogous behavior is found for the raw Fermi GBM dataset. The quadrupole amplitude remains systematically offset from the isotropic mean at the $\sim 3\sigma$ level for both $N_{\mathrm{side}}=32$ and $N_{\mathrm{side}}=64$ (Figures 24 and 25), while the dipole amplitude shows no statistically significant deviation from isotropy at any resolution (Figures 22 and 23). These results demonstrate that the quadrupole excess identified in the raw Fermi sky is likewise robust to reasonable variations in HEALPix resolution.
Taken together, these stability tests confirm that the moderate-significance quadrupole signals observed in the raw BATSE and Fermi GRB datasets are not artifacts of pixelization. While the exposure-corrected BATSE analysis shows that instrumental effects can fully account for this signal, particularly in that case, the persistence of the quadrupole excess across resolutions reinforces the need for careful exposure modeling particularly for Fermi GBM before definitive physical interpretations can be drawn.
Figure 18: Comparison of the observed dipole amplitude for the raw BATSE dataset (vertical downward arrow) with the estimated PDF obtained from Monte Carlo simulations of isotropic skies containing 2702 points each. The position of the observed dipole amplitude lies within the isotropic distribution and shows no statistically significant deviation. In this comparison, HEALPix resolution parameter $N_{\mathrm{side}}=32$ has been used.
Figure 19: Same as Figure 18, but with HEALPix resolution parameter $N_{\mathrm{side}}=64$. The observed dipole amplitude remains consistent with statistical isotropy, indicating stability against changes in pixelization scale.
Figure 20: Comparison of the observed quadrupole amplitude for the raw BATSE dataset (vertical downward arrow) with the estimated PDF obtained from Monte Carlo simulations of isotropic skies containing 2702 points each. The observed quadrupole amplitude remains elevated relative to the isotropic mean at a significance close to the $\sim 3.9\sigma$ level. In this comparison, HEALPix resolution parameter $N_{\mathrm{side}}=32$ has been used.
Figure 21: Same as Figure 20, but with HEALPix resolution parameter $N_{\mathrm{side}}=64$. The persistence of the elevated quadrupole amplitude confirms that the signal is not an artifact of the chosen pixelization scale.
Figure 22: Comparison of the observed dipole amplitude for the raw Fermi GBM dataset (vertical downward arrow) with the estimated PDF obtained from Monte Carlo simulations of isotropic skies containing 4032 points each. The observed dipole amplitude lies well within the isotropic distribution. In this comparison, HEALPix resolution parameter $N_{\mathrm{side}}=32$ has been used.
Figure 23: Same as Figure 22, but with HEALPix resolution parameter $N_{\mathrm{side}}=64$. The dipole amplitude remains statistically consistent with isotropy, demonstrating robustness to changes in HEALPix resolution.
Figure 24: Comparison of the observed quadrupole amplitude for the raw Fermi GBM dataset (vertical downward arrow) with the estimated PDF obtained from Monte Carlo simulations of isotropic skies containing 4032 points each. The observed quadrupole amplitude shows a deviation at the $\sim 3\sigma$ level relative to the isotropic mean. In this comparison, HEALPix resolution parameter $N_{\mathrm{side}}=32$ has been used.
Figure 25: Same as Figure 24, but with HEALPix resolution parameter $N_{\mathrm{side}}=64$. The quadrupole excess persists across resolutions, indicating that it is not driven by the choice of pixelization scale.